OpenStudy (anonymous):

MEDAL AWARD Help !! Find the equation of a line in slope-intercept form that passes through the points (-2, 20) and (3, -10).

jimthompson5910 (jim_thompson5910):

first you need the slope

jimthompson5910 (jim_thompson5910):

m = (y2-y1)/(x2-x1) m = ( -10- 20)/(3-(-2)) m = ( -10- 20)/(3+2) m = (-30)/5 m = -6 So the slope is m = -6

OpenStudy (anonymous):

your supposed to find the equation ?

jimthompson5910 (jim_thompson5910):

I know, you use that slope to do so

OpenStudy (anonymous):

how do u do that ?

jimthompson5910 (jim_thompson5910):

by using y = mx+b

jimthompson5910 (jim_thompson5910):

plug in m and (x,y) = (-2,20)

jimthompson5910 (jim_thompson5910):

*plug in m = -6

OpenStudy (anonymous):

-6(3,-10)=(-2,20)

jimthompson5910 (jim_thompson5910):

y = mx+b y = -6x+b 20 = -6(-2) + b solve for b

OpenStudy (anonymous):

i got 12 then you divide by 20 ??

jimthompson5910 (jim_thompson5910):

no

jimthompson5910 (jim_thompson5910):

20 = 12 + b 20 = b + 12 b = ??

OpenStudy (anonymous):

8 ..?

jimthompson5910 (jim_thompson5910):

correct

jimthompson5910 (jim_thompson5910):

so the equation is y = -6x + 8

OpenStudy (anonymous):

thanks

jimthompson5910 (jim_thompson5910):

np